A PARACHUTIST BAILS OUT AND FREELY FALLS 50M THEN THE PARACHUTE OPENS , AND THEREAFTER SHE DECELERATES AT 2M/S². SHE REACHES THE GROUND WITH A SPEED OF 3M/S.
1.) HOW LONG WAS THE PARACHUTE IN THE AIR?
2.) AT WHAT HEIGHT DID THE FALL BEGIN?
1
Expert's answer
2019-10-22T10:09:22-0400

There are two phases,first when she is without parachute for first 50 m and then with parachute.

First phase:"s_1=h=50 m;u=0;"

Let "v_1" be the velocity and "t_1" be the time elapsed after "50 m,so\ v_1=\sqrt{2gh}=\sqrt{2\times 9.8\times 50}=31.30\ m/sec"

"t_1=\sqrt{\frac{2h}{g}}=\sqrt{\frac{2\times 50}{9.8}}=3.194\ sec"

During second phase:"t_2"  be the time taken,Initial velocity"=v_1=31.30 \ m/sec;a=-2m/sec^2"

Final velocity"(v_2)=3m\/sec"

"\implies v_2=v_1+at_2\implies 3=31.30-2\times t_2\implies t_2=14.15 \ sec"

Distance traveled with parachute or height at which parachute open"=s_2=v_1t_2+\frac{1}{2}at_2^2;"

"s_2=31.30\times 14.15+\frac{1}{2}\times (-2)\times 14.15^2=" "242.67 m"

Hence "(a)" Total time parachutist was in air"=t_1+t_2=17.344 \ sec"

"(b)" The height at which fall begin"=s_1+s_2=292.67\ m"